Determine the value of EIS at the right end of the overhanging beam in fig. P-693.

900 N M = 600 Nm 2 m 3 m 2 m R1 R.2 2 Figura P-693.

Ans. EI\delta =680 N\cdot m^{3}
Answer #1

Page No 1 900 N M=600 Nm A 2m B D zm 2m * R R2 Using equilibrium equations, EMA= 0 → 900x2 -R2 (2+3)+600=0 5R2 = 2400 R2 = 24Using double integration method, Page No.2 El d2y = -Me >-El d2y = Max dx² dx² dxc² = +Rix - EI dy dx 2 2 - El d²Y 900 (4-2)Page. No. 3 At oc=Tm Y= 8 (deflection at point D - ET 8 = +420 x73 + 4.7+63 – 900 (+213 +480 (7-5) 6 6 = 5900 + 79 +2 Put val